what should we expect to be the result of 30e^4t=6
Answer #1

30e^4t = 6

=> e^4t = 6/30

=> e^4t = 1/5

Taking log of both sides,

=> 4t = log (1/5)

=> t = 4 × ( -1.609) = (-6.436)